OpenStudy (anonymous):

A plumber is fitting pipes for a remodeled kitchen sink. The sink requires 3 different sections of pipe; the first is 6√96 feet long, the second section is 12√150 feet long, and the third section is 2√294 feet long. How many feet of piping is required for all 3 sections?

OpenStudy (anonymous):

\[6 \sqrt{96}+12 \sqrt{150}+2 \sqrt{294}=98 \sqrt{6} \]

OpenStudy (anonymous):

@robtobey thank you so much :)